# Spectrum radii of trees

### v1

@@ 3 Weighted Directed trees
    for each $v\in V(T)$, $od_{T}(v)\leq r$ if $q=0$ and $v=z$, and $od_{T}(v)\leq r-1$ otherwise.
For example, the directed tree in Figure 2 belongs to both ${\mathscr{T}}_{3,2}(1)$ and ${\mathscr{T}}_{3,\sqrt{3}}(0)$. Two directed trees of ${\mathscr{T}}_{3,2}(\frac{1}{2})$ are also shown in Figure 3.
@@
    (a)                                                                  (b)
Figure 3: Two directed trees in $\mathcal{J}_{3,2}(\frac{1}{2})$
@@
Let $\alpha\geq 1$, $q\geq 0$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(q)$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$, unless $q=0$ and $v$ is the unique source of $T$.
Proof. Since $T\in{\mathscr{T}}_{r,\alpha}(q)$, we have $T\in{\mathscr{T}}_{\alpha}$. In the definition of ${\mathscr{T}}_{\alpha}$, the condition at (8) for the mapping $\omega_{\alpha,T}$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$ implies that for any vertex $v$ in $T$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ holds, unless $v$ is the source of $T$ and $q=0$. Hence the result follows. $\Box$
4 Proof of Theorem 1.1
Lemma 4.1.
Let $\alpha\geq 1$ and $r$ be a positive integer. If there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$.
Proof. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and (8), it is clear that ${\mathscr{T}}_{r,\alpha}(q)\neq\emptyset$ for any $q\in{\mathscr{W}}_{r}(\alpha)$.
Assume that $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. For each $i\in\llbracket s\rrbracket$, let $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, $\omega_{i}$ be the mapping $\omega_{\alpha,T_{i}}$ and $z_{i}$ be the unique source of $T_{i}$. By definition, $\omega_{i}(z_{i})=q_{i}$ for all $i\in\llbracket s\rrbracket$.
Now let $T$ be the directed tree obtained from $T_{1},T_{2},\dots,T_{s}$ by adding a new vertex $z$ and adding a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$. Let $\omega:V(T)\mapsto{\mathbb{R}}^{+}\cup\{0\}$ be the mapping defined by $\omega(z)=0$ and $\omega(w)=\omega_{i}(w)$ whenever $w\in V(T_{i})$. Since $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, condition (8) holds for $\omega$ at $z$. For any $v\in V(T)$, if $v\in V(T_{i})$, as $\omega(w)=\omega_{i}(w)$ for all $w\in V(T_{i})$, condition (8) also holds $\omega$ at $v$.
Since $s\leq r$ and $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket s\rrbracket$, we have $\Delta(T)\leq r$, we have $T\in{\mathscr{T}}_{r,\alpha}(0)$. $\Box$
Applying the set ${\mathscr{T}}_{r,\alpha}(0)$, we can now obtain the following conclusion for proving Theorem 1.1.
Lemma 4.2.
Let $\alpha\geq 1$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$. If $\alpha=1$, then $T\in{\mathscr{T}}_{r,\alpha}(0)$ implies that $|V(T)|=2$, and the result holds. Now assume that $\alpha>1$.
Let $T\in{\mathscr{T}}_{r,\alpha}(0)$. By definition, $T$ is a directed tree with a unique source, say $z$. By Lemma 3.1, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$. Clearly, $1\leq s\leq r$ and $T_{i}\in{\mathscr{T}}_{\alpha}$ for each $i\in\llbracket s\rrbracket$. Let $z_{i}$ be the vertex in $T_{i}$ such that $(z,z_{i})$ is a directed edge in $T$. Then, $z_{i}$ the unique source of $T_{i}$ and $q_{i}=\omega_{\alpha,T}(z_{i})$ for each $i\in\llbracket s\rrbracket$. By (8), we have
$0=\omega_{\alpha,T}(z)=\alpha-\sum_{i=1}^{s}q_{i}^{-1}.$ (9)
In order to prove that $\lambda(T)=\alpha$, by Proposition 2.3, it suffices to show that the mapping $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ defined below satisfies condition (4):
@@
    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$, where $\omega$ simply denotes $\omega_{\alpha,T}$.
We can first show that condition (4) is satisfied for the unique source and all sinks of $T$. As $z$ is the only source of $T$ and $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, where $N^{+}_{T}(z)$ is the set of out-neighbors of $z$ in $T$, we have
$\sum_{v\in N_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}q_{i}^{-1}=\alpha\phi(z),$ (10)
@@
For any sink $w$ of $T$, $\omega(w)=\alpha$ by (8). If $w^{\prime}$ is the unique in-neighbor of $w$ in $T$, then $\phi(w^{\prime})=\phi(w)\omega(w)$ by the definition of $\phi$, and
$\sum_{v\in N_{T}(w)}\phi(v)=\phi(w^{\prime})=\phi(w)\omega(w)=\alpha\phi(w).$ (11)
It remains to consider any vertex $w$ in $T$ which is neither the source nor a sink of $T$. Assume that $w^{\prime}$ ie the only in-neighbor of $w$ and $N^{+}_{T}(w)=\{w_{1},w_{2},\dots,w_{t}\}$. By the definition of $\phi$,
$\phi(w^{\prime})+\sum_{i=1}^{t}\phi(w_{i})=\omega(w)\phi(w)+\sum_{i=1}^{t}\frac{\phi(w)}{\omega(w_{i})}=\phi(w)\left(\omega(w)+\sum_{i=1}^{t}\omega(w_{i})^{-1}\right)=\alpha\phi(w),$ (12)
where the last step follows from (8). Thus, $\phi$ satisfies condition (4) for all vertices in $T$, implying that $\lambda(T)=\alpha$ by Proposition 2.3. $\Box$
Now we are going to establish the following conclusion, by which Theorem 1.1 follows directly.
Proposition 4.3.
For any positive numbers $\alpha$ and $r$, where $\alpha\geq 1$ and $r$ is an integer, the following three statements are pairwise equivalent:
  1. (i).
    there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$;
  2. (ii).
    $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$; and
  3. (iii).
    there exist $q_{1},q_{2},\dots,q_{s}\in{\mathscr{W}}_{r}(\alpha)$, where $1\leq s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$.
Proof. (iii) follows from (ii) directly. By Lemmas 4.2 and 4.1, (i) follows from (iii).
(i) $\Rightarrow$ (ii) Assume that $T$ is a tree with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$. By Proposition 2.3, there exists a mapping $\phi$ from $V(T)$ to ${\mathbb{R}}^{+}$ such that $\alpha\phi(u)=\sum\limits_{w\in N_{T}(u)}\phi(w)$ for each $u\in V(T)$. Then, by Lemma 2.4, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. Thus, (i) implies (ii). $\Box$
Clearly, Theorem 1.1 follows from Proposition 4.3 directly. We end this section by a result on the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
For any directed graph $D$, let $\Delta^{(o)}(D)$ denote the maximum out-degree of $D$. By Lemma 4.2 and its proof, the following conclusion holds.
Proposition 4.4.
Let $\alpha\geq 1$ and $r$ be an integer with $\alpha^{2}/4+1<r\leq\alpha^{2}$. Assume that there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$. If $s=r$ or $\Delta^{(o)}(T_{i})=r-1$ for some $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, where $i\in\llbracket s\rrbracket$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
@@
For any positive real number $\alpha$ and positive integer $r$ with $\alpha^{2}/4+1<r\leq\alpha^{2}$, if there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket r\rrbracket\}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
@@
Let $k$ be an integer with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. Clearly, for any integer $r$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$. If $k=4m+t$, where $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
In order to prove Theorem 1.4, by Corollary 4.5, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{r}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$. We start with the following simple result.
@@
    When $m=0$ and $1\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[1,t]$; and
@@
Thus, it mains to show that $\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ for $t=3$, i.e., $\frac{1}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$. Note that $\frac{2}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$. Thus,
$\frac{\sqrt{3}}{2}=\sqrt{3}-\frac{1}{\frac{2}{\sqrt{3}}}\in{\mathscr{W}}_{m+2}(\sqrt{3}),$
and
$\frac{1}{\sqrt{3}}=\sqrt{3}-\frac{1}{\frac{\sqrt{3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{3}).$
(ii) Assume that $m\geq 1$. For any integer $s$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
@@
We are now going to establish the conclusion that for any integer $k=4m+t$, where $m\geq 1$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ holds for all integers $s$ in the interval $[m,4m+t]$, on a case-by-case basis of $t$.
@@
For any positive integer $m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
@@
Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i\geq 1$.
By definition, $2\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus, the claim holds for $i=1$. For any positive integer $i$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
@@
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any integer $i$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
@@
Claim 4: For any integer $s$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Lemma 5.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any integer $i$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
@@
For any integer $i$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
@@
For any positive integer $m$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
@@
Claim 1: For all $i\geq 1$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
@@
Claim 2: For any integer $i$ with $1\leq i\leq m$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
@@
Claim 3: For any integer $i$ with $0\leq i\leq m$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
@@
Hence the claim holds for $i=0,1$. It suffices to show that for any integer $i$ with $20\leq i\leq m-$ 2, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Assume that $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then
@@
For any positive integer $m$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
@@
Claim 1: For any integer $i\geq 1$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
@@
It suffices to show that for any $i$ with $10\leq i\leq m-$ 1, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Observe that for any $i$ with $10\leq i\leq m-$ 1,
@@
By Claim 2, $\frac{2m+i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for $i=0,1$. For any integer $i$ with $20\leq i\leq m-$ 2, it is shown in (35) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
@@
For any positive integer $m$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.

### v2

@@ 1 Introduction
    $\omega(v)>0$ whenever $v\in V(T)\setminus\{z\}$, and
    $\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }N^{+}_{T}(v)=\emptyset;\\ \alpha-\sum\limits_{u\in N^{+}_{T}(v)}\omega(u)^{-1},&\mbox{otherwise},\end{array}\right.$ (5)
where $N^{+}_{T}(v)$ is the set of out-neighbors of $v$ in $T$ and $od_{T}(v)=|N^{+}_{T}(v)|$.
For any $\beta\in{\mathbb{R}}^{+}$, let ${\mathscr{T}}_{r,\alpha}(\beta)$ denote the set of ordered pairs $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}$ such that $\omega(z)=\beta$, where $z$ is the unique source of $T$. By the definitions of ${\mathscr{W}}_{r}(\alpha)$ and ${\mathscr{T}}_{r,\alpha}(\beta)$, the following conclusion holds.
Lemma 3.1.
For any $\beta\in{\mathscr{W}}_{r}(\alpha)$, ${\mathscr{T}}_{r,\alpha}(\beta)\neq\emptyset$.
Proof. The conclusion is trivial if $\beta=\alpha$. Now assume that $\beta=\alpha-\sum\limits_{i=1}^{s}\beta_{i}^{-1}$ and $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$ for each $i\in\llbracket s\rrbracket$, where $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$. Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{s}$ by adding a new vertex $z$ and a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$, where $z_{i}$ is the unique source of $T_{i}$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ such that $\omega(v)=\omega_{i}(v)$ holds for each $i\in\llbracket s\rrbracket$ and $v\in V(T_{i})$, and
$\omega(z)=\alpha-\sum\limits_{i=1}^{s}\omega(z_{i})^{-1}=\alpha-\sum\limits_{i=1}^{s}\beta_{i}^{-1}=\beta.$ (6)
Obviously, $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(\beta)$. Hence the lemma holds. $\Box$
For example, two members $(T_{i},\omega_{i})$ for $i=1,2$ of $\mathcal{J}_{3,2}(\frac{1}{2})$ are shown in Figure 2, where the value of $\omega_{i}(v)$ for each vertex $v$ in $T_{i}$ is shown beside vertex $v$.
@@
    (a) $(T_{1},\omega_{1})$                                             (b) $(T_{2},\omega_{2})$
Figure 2: $(T_{i},\omega_{i})\in\mathcal{J}_{3,2}(\frac{1}{2})$ for $i=1,2$
By Lemma 3.1, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
@@
For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$.
By definition again, $T$ is a directed tree with a unique source $z$ and $\omega(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$, where $s=od_{T}(z)$. Clearly, $1\leq s\leq r$. For each $i\in\llbracket s\rrbracket$, $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$, where $\omega_{i}$ is the restriction of $\omega$ to $V(T_{i})$, $\beta_{i}=\omega(z_{i})$ and $z_{i}$ is the vertex in $T_{i}$ with $z_{i}\in N^{+}_{T}(z)$. Then $z_{i}$ the unique source of $T_{i}$ for each $i\in\llbracket s\rrbracket$ and
$0=\omega(z)=\alpha-\sum_{i=1}^{s}\omega(z_{i})^{-1}=\alpha-\sum_{i=1}^{s}\beta_{i}^{-1}.$ (7)
Now let $\phi$ be the mapping from $V(T)$ to ${\mathbb{R}}^{+}$ defined below:
@@
    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$.
In order to show that $\lambda(T)=\alpha$, by Proposition 2.2, it suffices to prove the following Claim.
Claim A: $\phi$ satisfies the condition of (2) for each vertex $u\in V(T)$.
We first prove Claim A for the case $u=z$ or $u$ is a sink of $T$ (i.e., $N^{+}_{T}(u)=\emptyset$). As $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, we have
$\sum_{v\in N^{+}_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}\beta_{i}^{-1}=\alpha\phi(z),$ (8)
@@
For the case that $u$ is a sink of $T$, $\omega(u)=\alpha$ by (5). If $u^{\prime}$ is the unique in-neighbor of $u$ in $T$, then $\phi(u^{\prime})=\phi(u)\omega(u)$ by the definition of $\phi$, and
$\sum_{v\in N_{T}(u)}\phi(v)=\phi(u^{\prime})=\phi(u)\omega(u)=\alpha\phi(u).$ (9)
Finally, we assume that $u$ is vertex in $T$ which is neither the source nor a sink of $T$. Assume that $u^{\prime}$ ie the only in-neighbor of $u$ in $T$ and $N^{+}_{T}(w)=\{u_{1},u_{2},\dots,u_{t}\}$. By the definition of $\phi$,
$\phi(u^{\prime})+\sum_{i=1}^{t}\phi(u_{i})=\omega(u)\phi(u)+\sum_{i=1}^{t}\frac{\phi(u)}{\omega(u_{i})}=\phi(u)\left(\omega(u)+\sum_{i=1}^{t}\omega(u_{i})^{-1}\right)=\alpha\phi(u),$ (10)
where the last step follows from (5). Thus, $\phi$ satisfies condition (2) for each $u\in V(T)$, implying that $\lambda(T)=\alpha$ by Proposition 2.2. $\Box$
By Lemmas 3.1 and 3.2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$, and thus the sufficiency of Theorem 1.1 holds.
@@
If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
Proof. Assume that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, where $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ and $q_{i}>0$ for each $i\in\llbracket r\rrbracket$. By Lemma 3.1, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$. Assume that $z_{i}$ is the only source of $T_{i}$ for each $i\in\llbracket r\rrbracket$. Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{r}$ by adding a new vertex $z$ and directed edges $(z,z_{i})$ for each $i\in\llbracket r\rrbracket$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ defined as follows: $\omega(z)=0$ and $\omega(v)=\omega_{i}(v)$ whenever $v\in V(T_{i})$, where $i\in\llbracket r\rrbracket$. By the definition of $T$ and $\omega$, it can be verified easily that $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$. Clearly, $r\geq\Delta(T)\geq od_{T}(z)=r$, implying that $\Delta(T)=r$. By Lemma 3.2, $\lambda(T)=\alpha$. $\Box$
Let ${\mathscr{W}}_{r,\alpha}(0)=\big{\{}\omega(v):v\in V(T),(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)\big{\}}$.
Corollary 3.4.
Assume that $0\in{\mathscr{W}}_{r}(\alpha)$. If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
@@
Let $k\in{\mathbb{N}}^{+}$ with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. Clearly, for any $r\in{\mathbb{N}}^{+}$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$. If $k=4m+t$, where $m\in\{0\}\cup{\mathbb{N}}^{+}$ and $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
In order to prove Theorem 1.5, by Corollary 3.3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$. We start with the following simple result.
@@
    When $m=0$ and $2\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[2,t]$; and
@@
(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
@@
We are now going to establish the conclusion that for any $k=4m+t$, where $m\in{\mathbb{N}}^{+}$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+t})$ holds for all integers $s$ in the interval $[m,4m+t]$ by considering each possible value of $t$ separately.
@@
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
@@
Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i\in{\mathbb{N}}^{+}$.
By definition, $2\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus, the claim holds for $i=1$. For any $i\in{\mathbb{N}}^{+}$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
@@
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any $i\in{\mathbb{N}}^{+}$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
@@
Claim 4: For $s\in{\mathbb{N}}^{+}$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Lemma 4.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
@@
For any $i\in{\mathbb{N}}^{+}$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
@@
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
@@
Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
@@
Claim 2: For any $i\in\llbracket m\rrbracket$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
@@
Claim 3: For any $i\in\{0\}\cup\llbracket m\rrbracket$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
@@
Hence the claim holds for $i=0,1$. It suffices to show that for any $i\in\{0\}\cup\llbracket m-2\rrbracket$, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Assume that $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then
@@
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
@@
Claim 1: For any $i\in{\mathbb{N}}^{+}$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
@@
It suffices to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Observe that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$,
@@
By Claim 2, $\frac{2m+i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for $i=0,1$. For any $i\in\{0\}\cup\llbracket m-2\rrbracket$, it is shown in (33) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
@@
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.